OpenStudy (anonymous):

A store mixes Kenyan coffee worth $10 per kilogram and Turkish coffee worth $13 per kilogram. The mixture is to sell for $12 per kilogram. Find how much of each should be used to make a 492- kilogram mixture.

hero (hero):

10x + 13y = 492*12 x + y = 492

OpenStudy (anonymous):

so x = 6355/3 and and y = -4879/3 ?

hero (hero):

I get x = 164, y = 228

OpenStudy (anonymous):

thank you!